Using matrix method, solve the simultaneous equations {x-3y=3

5x-9y=11


1
Expert's answer
2021-09-20T17:34:28-0400

"A\bar x=\bar b"

"A=\begin{pmatrix}
 1 & -3 \\
 5 & -9
\end{pmatrix},\bar b=\begin{pmatrix}
 3 \\
 11
\end{pmatrix}" .

"det(A)=-9+15=6."

"A^{-1}=\frac{1}{6}\begin{pmatrix}
 -9 & 3 \\
 -5 & 1
\end{pmatrix}"

"\bar x=\begin{pmatrix}
 x \\
 y
\end{pmatrix}=A^{-1}\bar b=\frac{1}{6}\begin{pmatrix}
 -9& 3 \\
 -5 & 1
\end{pmatrix}\begin{pmatrix}
 3 \\
 11
\end{pmatrix}=\frac{1}{6}\begin{pmatrix}
 6 \\
 -4
\end{pmatrix}=\begin{pmatrix}
 1 \\
 -\frac{2}{3}
\end{pmatrix}"

Thus, "x=1,y=-\frac{2}{3}."